OpenStudy (anonymous):

Find x when y = 5 in the literal equation 4x + 7y = 51.

OpenStudy (anonymous):

4x + 7y = 51. 4x + 7(5) = 51 4x + 35 = 51 4x = 51 - 35 4x = 16 x = 16/4 = 4

OpenStudy (anonymous):

THANXX

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

can u help me with something els

OpenStudy (anonymous):

Valerie is ordering salads and drinks for her friends. Salads cost $7 each, drinks cost $3 each, and there is a $5 delivery charge per order. She has $50. If she buys 3 salads, how many drinks can she buy? If x = number of salads and y = number of drinks, which of the following inequalities could you use to answer the problem?